# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v2

@@ Abstract
Large language model (LLM) and agent techniques for data analysis (a.k.a LLM/Agent-as-Data-Analyst) have demonstrated substantial impact in both academica and industry. In comparison with traditional rule or small-model based approaches, (agentic) LLMs enable complex data understanding, natural language interfaces, semantic analysis functions, and autonomous pipeline orchestration. From a modality perspective, we review LLM-based techniques for $(i)$ structured data (e.g., table question answering for relational data and NL2GQL for graph data), $(ii)$ semi-structured data (e.g., markup languages understanding and semi-structured table modeling), $(iii)$ unstructured data (e.g., chart understanding, document understanding, code completion), and $(iv)$ heterogeneous data (e.g., data retrieval and modality alignment for data lakes). The technical evolution further distills five key design goals for intelligent data analysis agents, namely semantic-aware design, modality-hybrid integration, autonomous pipelines, tool-augmented workflows, and support for open-world tasks. Finally, we outline the remaining challenges and propose several insights and practical directions for advancing LLM/Agent-powered data analysis.
@@ 1 INTRODUCTION
Data analysis, broadly defined as the process of inspecting, transforming, and modeling data to discover useful information and support decision-making, constitutes a cornerstone of modern scientific research and business intelligence hardy2004handbook ; temporal2022 ; LLM4DM ; chat2data2024 . It spans a wide spectrum of data modalities, from structured databases and semi-structured tables to unstructured documents and videos, underpinning critical applications across domains such as finance, healthcare, engineering, and social sciences nooralahzadeh2024explainable ; batko2022use ; cao2022ai .
@@ 1.1 Limitations of Conventional Data Analysis
Traditional data analysis pipelines, although effective in extracting information and statistical patterns, often demand substantial domain expertise, manual feature engineering, and the integration of multiple specialized tools xiao2024cellagent . These limitations become more pronounced as data grows in scale, complexity, and heterogeneity liang2024natnl2gql , ultimately constituting the inherent weakness of conventional data analysis.
L1: Manual Development. The most labor-intensive challenge is that traditional data analysis workflows rely on rigid, manually designed pipelines that require domain experts to specify modeling steps in advance zhou2025cracksql ; zhou2025cracksqldemo ; zhang2023reactable . In large-scale enterprise databases, where separate tables for customers, orders, payments, shipments, and products each contain dozens of attributes, answering even simple business queries such as “What is the average delivery time for premium customers in the last quarter?” may necessitate the analysts to choose related tables, find corresponding attributes, process complex multi-table joins, apply domain-specific constraints, and reconcile the timestamp fields. For document query answering, retrieving query-relevant information from long documents containing diverse elements (e.g., text blocks, tables, charts, images) typically requires analysts to examine pages sequentially and summarize the content, which is a highly labor-intensive process. Such handcrafted pipelines are time-consuming to build and difficult to adapt when data characteristics or analysis goals change.
L2: Hard-Coded Tool Dependency.One observation is that existing data analysis systems often come with a suite of tightly coupled built-in tools (e.g., visualization dashboards, statistical packages, or machine learning modules torchptpth ) tailored to specific workflows, which stems from the inherent complexity of the underlying data types. For instance, to extract statistical information from a knowledge graph and subsequently model it with linear regression, a human analyst must first formulate GQL queries and invoke the built-in APIs of graph databases to obtain the data, and then employ machine learning libraries (e.g., scikit-learn, PyTorch) for further analysis. In contrast, analyzing 3D models relies more heavily on domain-specific expertise and specialized systems (e.g., AutoCAD, PyMOL pymol ). Such analyses are often constrained to the functionalities embedded in these systems, requiring analysts, particularly those without coding expertise, to integrate multiple systems in order to complete a single task. The integration of heterogeneous systems not only increases the complexity and overhead of the analysis process, but also hampers extensibility and complicates integration into broader analytical workflows.
@@
L3: Homogeneous Modality Support.Current data analysis systems are generally optimized for specific data types, with architectures and query engines tailored to particular formats. For example, relational databases are well-suited for structured data LLMQO , while systems such as MongoDB are designed for semi-structured formats like JSON mongodb . When analyzing videos with accompanying timestamped descriptions, human analysts must manually align frames with their corresponding documents to enable cross-modal analysis mu2024snagscalableaccuratevideo . This specialization hampers the analysis of heterogeneous data across modalities, thereby increasing analytical complexity, introducing errors, and limiting both the scope and efficiency of data-driven insights.
@@
L4: Basic Format-Based Analysis.While conventional analysis methods efficiently support operations such as filtering, aggregation, they lack the ability to reason over semantic information embedded in data. For example, in table analysis, if a cell contains lengthy textual paragraphs, SQL queries can only retrieve or match keywords without understanding the semantics (e.g., sentiment, intent, causal relationships) zhang2023reactable . Likewise, for unstructured data such as charts or documents, analysts must manually interpret and summarize content to fulfill analytical needs han2023chartllama . The absence of semantic understanding constrains the analytical depth of current systems and limits their ability to address complex, knowledge-intensive tasks that demand reasoning beyond surface-level patterns.
@@ 1.2 Opportunities by LLM/Agent-As-Data-Analyst
O1: Complex Data Understanding. LLMs are capable of processing and reasoning across complex data, including relational data, semi-structured tables, and unstructured texts, owing to their ability to capture latent patterns and contextual dependencies su2024tablegpt2 ; tang2024graphgpt . For example, they can not only comprehend the semantic content associated with nodes and edges but can also reason about the underlying structural properties (e.g., connectivity, community patterns, or hierarchical organizations). Such a deep and holistic understanding empowers analysts to derive richer insights that integrate quantitative measures with qualitative nuances arora2023language .
@@
O2: NL-Based Interface. An LLM-powered analysis agent leverages natural language as the primary interface for interaction, enabling users to articulate analytical requirements without the need for specialized query languages such as SQL or advanced programming skills liang2024natnl2gql ; birdsql . This natural language interface enhances convenience for analysts, accelerates the overall analysis process, and broadens participation in data-driven decision-making across diverse user groups tang2025st ; han2023chartllama .
O3: Semantic Operators. LLMs enable semantic-level operations such as structural interpretation and content summarization, which are particularly beneficial for data rich in structural relationships or textual content wang2024interactive . For example, when querying a medical record database, a conventional query may retrieve only documents containing the term “fever”, whereas an LLM-based operator can also recognize semantically related expressions such as “high temperature” or “febrile condition”. This semantic enrichment enhances the expressive power of queries, enabling analysts to obtain more accurate, context-aware results jiang2025explainable ; wang2024chainoftable .
O4: Autonomous Evolution. LLM agents function not as static tools but as adaptive systems capable of refining their performance through continuous interaction and learning. Such adaptability is essential, as analytical requirements and data environments inevitably evolve over time hong2024datainterpreter . For instance, a financial analysis agent may initially concentrate on trend forecasting but, through iterative feedback, expand its capabilities to detect anomalies in real time or incorporate emerging risk factors. This capacity for autonomous evolution ensures that the system remains advanced, alleviates the need for manual reconfiguration, and delivers increasingly personalized and efficient analytical support atreja2025alasautonomouslearningagent .
@@ 1.3 Techniques of LLM-Powered Data Analysis
$\bullet$ Document. Documents, such as PDFs, web pages, and scanned reports, serve as primary carriers of information in business and academia. Traditional document analysis relies on Optical Character Recognition (OCR) and rule-based template matching dococr ; ha2022information , which are often ineffective for complex or varied layouts and fail to grasp the deep semantics of the content. LLMs, particularly Multimodal LLMs, have revolutionized document understanding by unifying the processing of text, layout, and visual information in three aspects: $(i)$ Architectural innovations from the LayoutLM series to DocLLM, enable a synergistic understanding of document structure and content xu2020layoutlm ; xu2021layoutlmv2 ; huang2022layoutlmv3 ; wang2024docllm . Concurrently, $(ii)$ Retrieval-Augmented Generation (RAG) is employed for question answering and summarization across long or multiple documents lewis2020retrieval ; lei2024visdommultimodal , and $(iii)$ synthetic data generation effectively alleviates the scarcity of labeled data seol2024posterllama ; shi2025layoutcot . These techniques collectively drive the shift from simple information extraction to deep document reasoning and synthesis.
$\bullet$ Program. Program analysis, which aims to understand, verify, and optimize source code, is a cornerstone of software engineering. Traditional methods depend on static analysis (e.g., compiler theory) and dynamic analysis (e.g., unit testing), which, while rigorous, require extensive expertise and are difficult to scale to large, semantically complex codebases. By learning from vast amounts of code, LLMs have acquired powerful capabilities for code generation and comprehension, giving rise to new analytical paradigms. The core of this advancement lies in the construction and utilization of code-task pairs—paired data of code snippets with their specific attributes (e.g., vulnerability labels, functional descriptions). Through advanced data synthesis techniques like iterative refinement and self-correction luo2023wizardcoder ; chen2023teaching , models can generate high-quality training data, thereby empowering frontier applications such as Automated Theorem Proving (ATP), vulnerability detection, and repository-level code completion xin2024deepseek-prover ; liu2023gptvulnerability ; zhang2023repocoder . Notably, the deep evolution of Retrieval-Augmented Generation (RAG) in the code domain has significantly enhanced the model’s ability to leverage context from entire repositories liang2024repofuse ; wu2024repoformer ; guo2024ft2ra .
@@
LLM/Agent for Data Analysis Evolution. As shown in Figure 1, the technological evolution of LLM-powered Data Analysis Agents can be summarized along five key dimensions, each corresponding to the design goals of a unified data analysis agent.
$\bullet$ From Literal to Semantic. Early LLM-powered data analysis agents primarily assist data analysis by generating executable code or domain-specific languages (DSLs). Their evolution has shifted toward leveraging semantic understanding to derive analytical results directly, reducing the need for intermediary procedural steps.
$\bullet$ From Modality-Specific to Modality-Hybrid. Initial data analysis agents are limited to analyzing a single type of data. The current trend emphasizes the integration of multiple data modalities, enabling agents to perform coordinated analyses across heterogeneous data.
$\bullet$ From Manual Development to Autonomous Design. Early agents require manual decomposition of analytical workflows (e.g., task decomposition, code generation, operation execution). Modern agents increasingly support autonomous workflow design, with broader operational capabilities and reduced human intervention.
$\bullet$ From Tightly-Coupled to Flexible Tools. Traditional agents rely on tightly integrated, framework-specific tools. The evolution favors decoupled architectures, where LLMs can leverage arbitrary toolsets, enhancing flexibility and adaptability in diverse analytical contexts.
$\bullet$ From Closed World to Open World. Initially, agents are predominantly tailored to domain-specific tasks (e.g., financial or industry analysis). The trend is toward general-purpose agents capable of analyzing diverse and real-world data (e.g., documents, videos), lowering user entry barriers.
@@ 1.4 Comparison & Contributions
$\bullet$ Full-Vision Introduction to Typical Data Analysis Tasks. Different from existing surveys that typically focus on a single modality or task (e.g., NL2SQL sqlsurvey ; shi2024sqlsurvey ; zhu2024sqlsurvey , graph understanding shang2024graphsurvey ; ren2024graphsurvey , table question answering jin2022tablesurvey ; wu2025tablesurvey , document understanding ding2024docsurvey ; barboule2025docsurvey , chart understanding farahani2023chartsurvey ; al2024chartsurvey , video understanding tang2025videosurvey ), we systematically organize the landscape of data analysis by categorizing techniques across structured, semi-structured, unstructured, and heterogeneous data. This roadmap also traces the technical evolution of LLM-powered data analysis and identifies five key dimensions that serve as design goals for a general-purpose data analysis agent.
@@ 1.5 Organization of Our Survey
Section 2 discusses LLMs for structured data analysis, covering relational data (Section 2.1) and graph data (Section 2.2). Section 3 reviews LLMs for semi-structured data analysis, including markup languages (Section 3.1) and semi-structured tables (Section 3.2). Section 4 examines LLMs for unstructured data analysis, encompassing charts (Section 4.1), videos (Section 4.2), documents (Section 4.3), programming languages (Section 4.4), and 3D models (Section 4.5). Section 5 addresses LLMs for heterogeneous data analysis. For each data type, we first present the data analysis techniques, followed by a subsection on data curation. Finally, we discuss the challenges and future directions associated with each data type in Section 6, and conclude the survey in Section 7.
@@ 2 LLM for Structured Data Analysis
Structured data refers to data with well-defined schemas like relational data relationaldata and graph data graphdata . The commonality of them is pre-defined patterns that ensure clear organization and efficient querying. Relational data is characterized by row and column tables, or associated keys. Graph data models entities and their relationships through nodes and edges.
@@ 2.1 Relational Data Analysis
Figure 3: LLM for Structured Data Workflows - (a) Pipeline Method. (b) End-to-End Method.
$\bullet$ NL2SQL. With the help of LLM, users can directly perform operations using natural language. NL2SQL focuses on translating natural language queries into SQL commands by leveraging techniques such as $(i)$ schema linking, which aligns user intents with database schema to resolve ambiguities and filters out irrelevant schema to enhance the conciseness of input zhang2024finsql ; petsql , $(ii)$ content retrieval, which dynamically extracts relevant information from the database to refine query generation talaei2024chess ; li2024codes , and $(iii)$ SQL generation strategies such as multi-step generation that concentrates on different SQL parts at each step fan2024combining ; zhou2025cracksql ; zhou2025cracksqldemo , intermediate SQL representation or hybrid table representation for better LLM comprehension DINSQL ; DBLP:journals/pacmmod/XieXZG25 ; long2025bridging , and different decoding strategies (e.g., beam search, greedy search, or PICARD SQL360 ). Other surveys provide more comprehensive overviews for NL2SQL techniques Liu2024ASO ; luo2025natural ; shi2024survey .
$\bullet$ NL2Code. Different from NL2SQL, NL2Code approaches emphasize enhancing relational data analysis through generating Python code (e.g., Pandas, NumPy), which includes a vast number of library APIs characterized by high variability and complexity, and often requiring the handling of complex chain operations. Recent studies have made progress in alleviating these challenges, though many remain open.
@@
LLM for Semantic Analysis. Moreover, tasks requiring semantic understanding or natural language outputs (e.g., table summarization) benefit from LLM-based methods, including $(i)$ multi-step question answering with decomposition strategies and $(ii)$ end-to-end QA using optimized LLMs.
@@
(2) Iterative Decomposition: Static decomposition often struggles with multi-hop queries, whereas LLM-driven iterative decomposition dynamically overcomes it by refining sub-tasks via recursive reasoning. TAPERA zhao2024tapera applies this approach by introducing query decomposition, where the Planner breaks the query into sub-queries, the Reasoner generates executable programs for each, and the Answer Generator derives results to fulfill the plan. Finally, the Planner updates or finalizes the plan as needed. Similarly, ReAcTable zhang2023reactable and CHAIN-OF-TABLE wang2024chainoftable iteratively generate operations and update tables, using LLMs and in-context learning to produce reasoning chains as proxies for intermediate thoughts.
$\bullet$ End-to-End QA. End-to-End Question Answering refers to methods where the LLM directly generates the final answer without intermediate steps or iterative refinement. Approaches are classified by data handling into table-specific LLM fine-tuning, table content retrieval, and table-as-image analysis.
@@
(2) Table Content Retrieval: Rather than embedding entire tables, content retrieval selects relevant rows and columns to reduce noise. CABINET patnaik2024cabinet employs a weakly supervised component to produce a parsing statement that defines the criteria for selecting relevant cells, while TableMaster cao2025tablemaster refines subtables via row and column lookup, ranks candidate columns using LLM prompts, selects query-relevant subsets, and generates SQL queries to extract the most pertinent rows. On the Multi-table and Multi-hop QA dataset (MMQA) wu2025mmqa , LLMs must retrieve relevant tables from a corpus and perform multiple inferences. Iterative multi-table retrieval (MTR), fetching question- and table-related tables stepwise, achieves the best performance.
(3) Table-As-Image Analysis: To overcome text-only LLM limitations in understanding table structures, tables are converted to images for multimodal LLM analysis. Table-LLaVA zheng2024multimodal pretrains LLaVA-7B liu2024llava on 150K table recognition samples (e.g., output HTML of the input table) to align table structures and elements with textual modality and fine-tunes on 232K downstream tasks (e.g., fact verification, question answering) to enhance its instruction-following ability. TabPedia zhao2024tabpedia enables a single model to handle diverse table analysis tasks via a concept-synergy mechanism that abstracts tasks into concepts. Built on Vicuna-7B zheng2023judging , it adds meditative tokens to the LLM decoder input, adaptively activating visual token regions to interpret task-specific intents. However, limitations remain for such methods when processing twisted, distorted tables and direct document-image inputs.
LLM for Time Series Analysis. Time series data, a specialized form of relational data stored in relational or time-series databases (e.g., InfluxDB naqvi2017time ), captures sequential observations over time, with each step recording one or more variables. Analysis focuses on temporal trends, periodic patterns, and inter-variable dependencies (e.g., causality, correlation, coupling) to support prediction, classification, and detection liu2025towards . Unlike static relational analysis, time series emphasizes the dynamic evolution of variables along the timeline.
Traditional time series analysis relies on statistical and machine-learning methods, such as ARIMA and LSTM 8614252 . However, statistical approaches are limited by linear assumptions, manual feature engineering, and difficulty capturing complex nonlinear patterns rahmani2022association . Advances in machine and deep learning offer more flexible solutions 9005997 , alongside efforts to improve result interpretability 9713986 . Recently, LLMs have emerged for time series analysis, leveraging their sequence modeling, generalization capacity, and ability to capture long-term dependencies.
$\bullet$ TS2NL. Due to the sequential similarity between time series and natural language, TS2NL converts time series data into natural language for LLMs, enhancing intuitiveness and interpretability.

### v3

# Chunwei Liu9, Bin Wang4, Conghui He4, Xiaoyang Wang6, Fan Wu1

@@ Abstract
Large language models (LLMs) and agent techniques have brought a fundamental shift in the functionality and development paradigm of data analysis tasks (a.k.a. LLM/Agent-as-Data-Analyst), demonstrating substantial impact across both academia and industry. In comparison with traditional rule or small-model based approaches, (agentic) LLMs enable complex data understanding, natural language interfaces, semantic analysis functions, and autonomous pipeline orchestration. From a modality perspective, we review LLM-based techniques for $(i)$ structured data (e.g., NL2SQL, NL2GQL, ModelQA), $(ii)$ semi-structured data (e.g., markup languages understanding, semi-structured table question answering), $(iii)$ unstructured data (e.g., chart understanding, text/image document understanding), and $(iv)$ heterogeneous data (e.g., data retrieval and modality alignment in data lakes). The technical evolution further distills four key design goals for intelligent data analysis agents, namely semantic-aware design, autonomous pipelines, tool-augmented workflows, and support for open-world tasks. Finally, we outline the remaining challenges and propose several insights and practical directions for advancing LLM/Agent-powered data analysis.
@@ 1 INTRODUCTION
Data analysis, broadly defined as the process of inspecting, transforming, and modeling data to discover useful information and support decision-making, constitutes a cornerstone of modern scientific research and business intelligence hardy2004handbook ; temporal2022 ; LLM4DM ; chat2data2024 . It spans a wide spectrum of data modalities, from structured databases and semi-structured tables to unstructured documents and videos, underpinning critical applications across domains such as finance cao2022ai , healthcare batko2022use , and social sciences nooralahzadeh2024explainable .
@@ 1.1 Limitations of Conventional Data Analysis
Traditional data analysis pipelines, although effective in extracting information and statistical patterns, often demand substantial domain expertise, manual feature engineering, and the integration of multiple specialized tools xiao2024cellagent . These limitations become more severe as data grows in scale, complexity, and heterogeneity liang2024natnl2gql , ultimately constituting the inherent weakness of conventional data analysis.
@@
L1: Closed-World Assumption. Current data analysis systems are generally optimized for specific data types, with architectures and query engines tailored to particular formats. For example, relational databases are well-suited for structured data LLMQO , while systems such as MongoDB mongodb are designed for semi-structured formats like JSON. When analyzing videos with accompanying timestamped descriptions, human analysts must manually align frames with their corresponding documents to enable cross-modal analysis mu2024snagscalableaccuratevideo . This specialization hampers the analysis of heterogeneous data across modalities, thereby increasing analytical complexity, introducing errors, and limiting both the scope and efficiency of data-driven insights.
L2: Rigid Tool Coupling. One observation is that existing data analysis systems often come with a suite of tightly coupled built-in tools (e.g., visualization dashboards, statistical packages, or machine learning modules torchptpth ) tailored to specific workflows, which stems from the inherent complexity of the underlying data types. For instance, to extract statistical information from a knowledge graph and subsequently model it with linear regression, a human analyst must first formulate GQL queries and invoke the built-in APIs of graph databases to obtain the data, and then employ machine learning libraries (e.g., scikit-learn, PyTorch) for further analysis. In contrast, analyzing 3D models relies more heavily on domain-specific expertise and specialized systems (e.g., AutoCAD, PyMOL pymol ). Such analyses are often constrained to the functionalities embedded in these systems, requiring analysts, particularly those without coding expertise, to integrate multiple systems in order to complete a single task. The integration of heterogeneous systems not only increases the complexity and overhead of the analysis process, but also hampers extensibility and complicates integration into broader analytical workflows.
@@
L3: Basic Literal Analysis. While conventional analysis methods effectively support operations such as filtering and aggregation, they operate primarily at the literal level and lack the ability to reason over the semantic content embedded within data. For example, in table analysis, when a cell contains lengthy textual descriptions, SQL queries can only retrieve or match keywords, without interpreting deeper meanings such as sentiment, intent, or causal relationships zhang2023reactable . Similarly, for unstructured data types like charts or documents, analysts are often required to manually interpret and summarize the content to meet analytical goals han2023chartllama . This reliance on surface-level processing limits the analytical depth of current systems and hinders their capacity to handle complex, knowledge-intensive tasks that require semantic reasoning, abstraction, or contextual understanding beyond mere syntactic patterns.
L4: Manual-Based Development. Manual efforts in data analysis pose significant bottlenecks in both workflow construction and system design.
(1) Static Analysis Workflow Development. A major source of labor intensity lies in the reliance on rigid, manually crafted analytical pipelines, which require domain experts to predefine modeling steps zhou2025cracksql ; zhou2025cracksqldemo ; zhang2023reactable . In large-scale enterprise databases, where data is distributed across separate tables, even a simple business question such as “What is the average delivery time for premium customers in the last quarter?” demands a series of complex manual operations. Analysts must identify the relevant tables, select appropriate attributes, compose multi-table joins, impose domain-specific constraints, and align timestamp semantics. Similarly, in document-centric analytics, retrieving query-relevant content from lengthy, heterogeneous reports (comprising text, tables, charts, and images) often involves sequentially reading through pages and manually synthesizing information, which is both time-consuming and error-prone.
(2) Manual System Design and Engineering. Beyond workflow construction, existing data analysis systems themselves are heavily dependent on manual engineering, from designing features and crafting rule sets to integrating models, resulting in high development overhead, limited scalability, and poor adaptability to new analysis tasks or data modalities. Such reliance on static, human-defined logic restricts the ability of systems to generalize across use cases and hinders progress toward autonomous, self-evolving analytical tools.
@@ 1.2 Opportunities by LLM/Agent-As-Data-Analyst
O1: Cross-Modal Open World Data Support. LLMs are capable of processing and reasoning across complex data, including relational data, semi-structured tables, and unstructured texts, owing to their ability to capture latent patterns and contextual dependencies su2024tablegpt2 ; tang2024graphgpt . For example, they cannot only comprehend the semantic content associated with nodes and edges but can also reason about the underlying structural properties (e.g., connectivity, community patterns, or hierarchical organizations). Such a deep and holistic understanding empowers analysts to derive richer insights that integrate quantitative measures with qualitative nuances arora2023language .
@@
O2: NL-based Tool Agnostic Interaction.LLM-powered analysis agents employ natural language as the primary interface for interaction, allowing users to express analytical intents without relying on specialized query languages such as SQL or advanced programming skills liang2024natnl2gql ; birdsql . By abstracting away tool-specific interfaces and technical barriers, this natural language interaction paradigm enhances usability, streamlines the analysis process, and significantly lowers the entry threshold for non-expert users. As a result, it broadens participation in data-driven decision-making and facilitates more inclusive access to complex analytical capabilities across diverse user groups tang2025st ; han2023chartllama .
O3: Semantic Operators. LLMs empower a new class of semantic-level operators, such as semantic aggregation, semantic filtering, and semantic joins, that go beyond syntactic matching to interpret and manipulate data based on latent meaning wang2024interactive . These operators are particularly beneficial when working with data that contains rich textual descriptions or complex relationships. For example, in a medical records database, a conventional query might aggregate all entries containing the term “fever”, whereas a semantic aggregation operator can group or filter entries based on synonymous or contextually related expressions such as “high temperature” or “febrile condition”. Similarly, semantic joins allow for linking records across tables even when the join keys are not exact matches but share implicit meaning (e.g., “Type II Diabetes” and “adult-onset diabetes”) jiang2025explainable ; wang2024chainoftable . By embedding these capabilities directly into the query execution layer, LLMs significantly enhance the expressiveness and flexibility of analytical workflows.
O4: Automatic Workflow Orchestration &\& Evolution. First, LLM-powered analysis agents offer the potential for autonomous orchestration of analytical workflows by dynamically selecting appropriate data sources, operations, and tools based on user intent and data context. Unlike traditional static pipelines, such agents can reason about task requirements and adaptively construct end-to-end workflows, reducing the burden on analysts to manually specify each step. Beyond orchestration, recent research also explores the possibility of autonomous system evolution, such as automatically designing prompts, selecting tools, or even generating new agentic modules tailored to specific tasks. While these capabilities remain underexplored, they represent a promising future direction toward self-improving, minimally supervised analytical systems.
@@ 1.3 Techniques of LLM-Powered Data Analysis
$\bullet$ Document. Documents, such as PDFs, web pages, and scanned reports, serve as primary carriers of information in business and academia. Traditional document analysis relies on Optical Character Recognition (OCR) and rule-based template matching dococr ; ha2022information , which are often ineffective for complex or varied layouts and fail to grasp the deep semantics of the content. LLMs, particularly Multimodal LLMs, have revolutionized document understanding by unifying the processing of text, layout, and visual information in three aspects: $(i)$ Architectural Design from the LayoutLM xu2020layoutlm ; xu2021layoutlmv2 ; huang2022layoutlmv3 series to DocLLM wang2024docllm , enable a synergistic understanding of document structure and content. Concurrently, $(ii)$ Visual-Based Understanding is employed for question answering and summarization across documents with diverse layouts and elements lewis2020retrieval ; lei2024visdommultimodal , and $(iii)$ Text-Based Understanding effectively solves the document understanding on long and text-rich documents with techniques like RAG lin2024zendb ; sun2025quest . These techniques collectively drive the shift from simple information extraction to deep document reasoning.
@@
LLM/Agent for Data Analysis Evolution. As shown in Figure 1, the technological evolution of LLM-powered Data Analysis Agents can be summarized along four key dimensions, each corresponding to the design goals of a unified data analysis agent.
$\bullet$ From Closed World to Open World. Initially, agents are predominantly tailored to domain-specific tasks and types(e.g., financial or industry analysis). The trend is toward general-purpose agents capable of analyzing diverse and real-world data (e.g., documents, videos), lowering user entry barriers.
$\bullet$ From Rigid Tool Coupling to Flexible Tools. Traditional agents rely on tightly integrated, framework-specific tools. The evolution favors decoupled architectures, where LLMs can leverage arbitrary toolsets, enhancing flexibility and adaptability in diverse analytical contexts.
$\bullet$ From Basic Literal Analysis to Semantic Supported Analysis. Conventional analysis methods primarily operate at the literal level and lack the ability to reason over the semantic content embedded within data. The evolution has shifted toward leveraging semantic reasoning, abstraction, or contextual understanding to derive analytical results.
$\bullet$ From Manual Development to Fully Autonomous Design. Early agents require manual decomposition of analytical workflows (e.g., task decomposition, code generation, operation execution). Modern agents increasingly support both autonomous analysis workflow design and analysis system design, with broader operational capabilities and reduced human intervention.
@@ 1.4 Comparison & Contributions
$\bullet$ Full-Vision Introduction to Typical Data Analysis Tasks. Different from existing surveys that typically focus on a single modality or task (e.g., NL2SQL sqlsurvey ; shi2024sqlsurvey ; zhu2024sqlsurvey , graph understanding shang2024graphsurvey ; ren2024graphsurvey , table-QA jin2022tablesurvey ; wu2025tablesurvey , document understanding ding2024docsurvey ; barboule2025docsurvey , chart understanding farahani2023chartsurvey ; al2024chartsurvey , video understanding tang2025videosurvey ), we systematically organize the landscape of data analysis by categorizing techniques across structured, semi-structured, unstructured, and heterogeneous data. This roadmap also traces the technical evolution of LLM-powered data analysis and identifies five key dimensions that serve as design goals for a general-purpose data analysis agent.
@@ 1.5 Organization of Our Survey
Section 2 discusses LLMs for structured data analysis, covering relational data (Section 2.1) and graph data (Section 2.2). Section 3 reviews LLMs for semi-structured data analysis, including markup languages (Section 3.1) and semi-structured tables (Section 3.2). Section 4 examines LLMs for unstructured data analysis, encompassing charts (Section 4.1), videos (Section 4.2), documents (Section 4.3), and 3D models (Section 4.4). Section 5 addresses LLMs for heterogeneous data analysis. For each data type, we first present the data analysis techniques, followed by a subsection on data curation. Section 6 introduces some industry practices of LLM-powered data analysis. Finally, we discuss the challenges and future directions associated with each data type in Section 7, and conclude the survey in Section 8.
@@ 2 LLM for Structured Data Analysis
Structured data refers to data with well-defined schemas like relational data relationaldata and graph data graphdata . Both of them share pre-defined patterns that enable clear organization and efficient querying. Relational data is characterized by row and column tables, or associated keys. Graph data models entities and their relationships through nodes and edges.
@@ 2.1 Relational Data Analysis
Figure 3: LLM for Structured Data Analysis - (a) Pipeline Method. (b) End-to-End Method.
$\bullet$ NL2SQL. NL2SQL focuses on translating natural language queries into SQL commands by leveraging techniques such as $(i)$ schema linking, which aligns user intents with database schema to resolve ambiguities and filters out irrelevant schema to enhance the conciseness of input zhang2024finsql ; petsql , $(ii)$ content retrieval, which dynamically extracts relevant information from the database to refine query generation talaei2024chess ; li2024codes , and $(iii)$ SQL generation strategies such as multi-step generation that concentrates on different SQL parts at each step fan2024combining ; zhou2025cracksql ; zhou2025cracksqldemo , intermediate SQL representation or hybrid table representation for better LLM’s comprehension DINSQL ; DBLP:journals/pacmmod/XieXZG25 ; long2025bridging , and different decoding strategies (e.g., beam search, greedy search, or PICARD SQL360 ). Other surveys provide more comprehensive overviews for NL2SQL techniques Liu2024ASO ; luo2025natural ; shi2024survey .
$\bullet$ NL2Code. Unlike NL2SQL, NL2Code approaches focus on enhancing relational data analysis by generating Python code (e.g., using Pandas or NumPy). These methods must cope with a vast number of library APIs that are highly variable and complex, often involving intricate chained operations. Recent studies have made progress in alleviating these challenges, though many remain open.
@@
LLM for Semantic Analysis. Moreover, tasks that demand semantic understanding or natural language generation (e.g., table summarization) also benefit from LLM-based methods, including $(i)$ multi-step question answering with decomposition strategies and $(ii)$ end-to-end QA using optimized LLMs.
@@
(2) Iterative Decomposition: Static decomposition often struggles with multi-hop queries, whereas LLM-driven iterative decomposition dynamically overcomes it by refining sub-tasks via recursive reasoning. TAPERA zhao2024tapera applies this approach by introducing query decomposition, where the Planner breaks the query into sub-queries, the Reasoner generates executable programs for each, and the Answer Generator derives results to fulfill the plan. Finally, the Planner updates or finalizes the plan as needed. Similarly, ReAcTable zhang2023reactable and CHAIN-OF TABLE wang2024chainoftable iteratively generate operations and update tables, using LLMs and in-context learning to produce reasoning chains as proxies for intermediate thoughts.
$\bullet$ End-to-End QA. End-to-End Question Answering refers to methods where the LLM directly generates the final answer without intermediate steps or iterative refinement. These Approaches are classified by data handling into $(i)$ table-specific LLM fine-tuning, $(ii)$ table content retrieval, and $(iii)$ table-as-image analysis.
@@
(2) Table-As-Image Analysis: To overcome the limitations of text-only LLM in understanding table structures, tables are converted to images for multimodal LLM analysis. Table-LLaVA zheng2024multimodal pretrains LLaVA-7B liu2024llava on 150K table recognition samples (e.g., output HTML of the input table) to align table structures and elements with textual modality and fine-tunes on 232K downstream tasks (e.g., fact verification, question answering) to enhance its instruction-following ability. TabPedia zhao2024tabpedia enables a single model to handle diverse table analysis tasks via a concept-synergy mechanism that abstracts tasks into concepts. Built on Vicuna-7B zheng2023judging , it adds meditative tokens to the LLM decoder input, adaptively activating visual token regions to interpret task-specific intents.
LLM for Time Series Analysis. Time series data, a specialized type of relational data stored in relational or time-series databases (e.g., InfluxDB naqvi2017time ), records sequential observations over time, each capturing one or more variables. Its analysis targets temporal trends, periodic patterns, and inter-variable dependencies (e.g., causality, correlation, coupling) to enable prediction, classification, and detection liu2025towards . Unlike static relational analysis, time series emphasizes the dynamic evolution of variables along the timeline.
Traditional time series analysis relies on statistical and machine-learning methods, such as ARIMA and LSTM 8614252 . However, statistical approaches are limited by linear assumptions, manual feature engineering, and difficulty capturing complex nonlinear patterns rahmani2022association . Advances in machine and deep learning offer more flexible solutions 9005997 , alongside efforts to improve result interpretability 9713986 . Recently, LLMs have been applied to time series analysis, leveraging their strengths in sequence modeling, generalization, and long-term dependency capture.
$\bullet$ TS2NL. Due to the sequential similarity between time series and natural language, TS2NL converts time series data into natural language to enhance intuitiveness and interpretability for LLMs.